Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.50 coming 3 years from today. The dividend should grow rapidly - at a rate of 28% per year - during Years 4 and 5, but after Year 5, growth should be a constant 6% per year. If the required return on Computech is 12%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Year 3 dividend = 1.5

Year 4 dividend = 1.5 (1 + 28%) = 1.92

Year 5 dividend = 1.92 (1 + 28%) = 2.4576

Year 6 dividend = 2.4576 (1 + 6%) = 2.605056

Value at year 5 = D6 / required rate - growth rate

Value at year 5 = 2.605056 / 0.12- 0.06

Value at year 5 = 2.605056 / 0.06

Value at year 5 = 43.4176

Value of the stock = 1.5 / (1 + 0.12)3 + 1.92 / (1 + 0.12)4 + 2.4576 / (1 + 0.12)5 + 43.4176 / (1 + 0.12)5

Value of the stock = $28.32